Grade 11General Physics

How to find percentage decrease in the weight of the body when take to height 32 km above the surface of earthif radius of earth is 6400 km...Ans is 1 % how?

Harpreet
9 Years agoGrade 11

2 Answers

Vishal
9 Years ago
g(h)= g(1-2h/r) { g(h)- gravity at height h % decrease= 1 Kitna likhu bhai? Samjh ja ab I am anonymous trying to help but these people are not allowing me to do so.
Sonia Paruthi
9 Years ago
Here d= 32 km R= 6400 kmweight of body at depth d is mg` =mg(1-d/R)% decrease in weight =( mg- mg`)/mg × 100= 32 / 6400×100 = 0.5%.